# RealHiTBench: A Comprehensive Realistic Hierarchical Table Benchmark for Evaluating LLM-Based Table Analysis

### v1

@@ B.3 QA prompt
We also introduce answer prompts to generate various questions according to different task type with GPT. Apart from the fundamental prompt parts, such as Role Play and Output Control, we apply COT (Wei et al., 2022) to the answer prompt to improve the capability of understanding. All answer prompts are presented from Table 13 to Table 16. Notably, because the answer content and formats obviously vary among different subtypes of Data Analysis, we design detailed answer prompt for each subtype (Here we just display the prompt of rudimentary analysis.).
@@ Appendix C More Implementation Details
QwQ-32B Text 28.95 43.13 17.94 33.57 12.99 37.96 68.66 20.86 31.25 12.5
Doubao-1.5-pro-32k Text 59.65 66.34 45.4 53.77 65.9 72.45 75.21 36.01 25.97 14.94
DeepSeek-R1-Distill-Qwen-1.5B Text 11.18 17.75 7.13 15.74 9.92 19.93 38.02 16.24 14.94 0
DeepSeek-R1-Distill-Qwen-7B Text 28.1 34.53 17.64 24.96 19.08 28.1 54.39 25.19 39.61 6.49
DeepSeek-R1-Distill-Llama-8B Text 22.43 33.29 15.18 24.64 14.76 26.66 54.55 26.98 12.99 5.19
mPLUG-Owl2-7B Image 3.46 7.23 1.15 4.72 9.14 14.21 25.46 12.13 1.53 0
Qwen2-VL-7B-Instruct Image 28.5 38.34 9.34 22.88 45.29 55.99 37.41 24.3 16.23 5.84
Qwen2-VL-7B-Instruct Image+Text 32.79 45.19 11.54 24.97 48.6 60.42 39.69 25.39 25.32 9.09
@@ E.1 Complex Tables Examples
[Question1]: … ,
[subtype1]: … .
[Question2]: … ,
[subtype2]: … .
@@
Figure 13: The flat table sample in TableBench (Wu et al., 2024).
Figure 14: The multiple table sample in MMQA (Wu et al., 2025).

### v2

@@ B.3 QA prompt
We also introduce answer prompts to generate various questions according to different task type with GPT. Apart from the fundamental prompt parts, such as Role Play and Output Control, we apply COT (COT) to the answer prompt to improve the capability of understanding. All answer prompts are presented from Table 13 to Table 16. Notably, because the answer content and formats obviously vary among different subtypes of Data Analysis, we design detailed answer prompt for each subtype (Here we just display the prompt of rudimentary analysis.).
@@ Appendix C More Implementation Details
QwQ-32B Text 43.13 28.95 33.57 17.94 37.96 12.99 20.86 68.66 31.25 12.50
Doubao-1.5-pro-32k Text 66.34 59.65 53.77 45.40 72.45 65.90 36.01 75.21 25.97 14.94
DeepSeek-R1-Distill-Qwen-1.5B Text 17.75 11.18 15.74 7.13 19.93 9.92 16.24 38.02 14.94 0
DeepSeek-R1-Distill-Qwen-7B Text 34.53 28.10 24.96 17.64 28.10 19.08 25.19 54.39 39.61 6.49
DeepSeek-R1-Distill-Llama-8B Text 33.29 22.43 24.64 15.18 26.66 14.76 26.98 54.55 12.99 5.19
mPLUG-Owl2-7B Image 7.23 3.46 4.72 1.15 14.21 9.14 12.13 25.46 1.53 0
Qwen2-VL-7B-Instruct Image 38.34 28.50 22.88 9.34 55.99 45.29 24.30 37.41 16.23 5.84
Qwen2-VL-7B-Instruct Image+Text 45.19 32.79 24.97 11.54 60.42 48.60 25.39 39.69 25.32 9.09
@@ E.1 Complex Tables Examples
[Question1]: … , [subtype1]: … .
[Question2]: … , [subtype2]: … .
@@
Figure 13: The flat table sample in TableBench (tablebench).
Figure 14: The multiple table sample in MMQA (MMQA).